Which equation is true for the value x = 15?
2(x + 3) = 40
2(x − 5) = 30
2(x + 5) = 40
x + 2x = 30
3x − x = 45

Answer by stanbon(75887)

Which equation is true for the value x = 15?
2(x + 3) = 40
2(15+3) = 40
2*18 = 40
False
-----------
2(x − 5) = 30
2(15-5) = 30
2(10) = 30
False
----------------
2(x + 5) = 40
2(15+5) = 40
2(20) = 40
40 = 40
True
-------------------
x + 2x = 30
3x − x = 45
=============
Cheers,
Stan H.
----------